OpenStudy (anonymous):

Solve the following (leave answer in radical form). x^2-4x+5=0

OpenStudy (anonymous):

x = [ 4 +/- sqrt (-4) ] / (2)

OpenStudy (anonymous):

x=2+/-i

OpenStudy (anonymous):

x= 2+i x=2-i